# Accelerating shape optimization by deep neural networks with on-the-fly determined architecture

### v1

@@ 4 Benchmarking of CFDNNetAdapt
Figure 3: Comparison of solutions to the ZDT problems. From left to right, there are optimal solutions (black), solutions found by NSGA-II (red), solutions found by CFDNNetAdapt (blue), solutions found by SOCEMO (orange), and solutions found by a CFDNNetAdapt variant with GPRs (green). The empty circles depict the Pareto-optimal front as predicted by the surrogate model, the full circles show the corresponding true values.
@@
The solutions (Pareto-optimal fronts) found by NSGA-II are shown in Figure 3 in red. Similarly to deb2002, NSGA-II was run with population size $n_{\mathrm{pop}}=250$ and number of generations $n_{\mathrm{gen}}=100$. All evaluated data points (tuples ($\bm{p}$$\bm{o}$)) generated by NSGA-II ($\sim\!\!25,000$) were used to feed the other SAOAs.
CFDNNetAdapt solutions are depicted in blue. For each ZDT problem, two Pareto-optimal fronts corresponding to the same Pareto-optimal set are depicted: the one predicted by the best found surrogate model (empty circles) and its counter-part re-evaluated by the true cost functions (full circles). The results of the other SAOAs are depicted in the same way. Moreover, for all optimization problems, the same hyperparameter setting was used to keep the comparison fair; the setting is discussed in C.
The solutions of SOCEMO [muller2017] are shown in orange. The surrogates used by SOCEMO are based on radial basis functions (RBFs) and were originally used for problems with up to $500$ data points. As we work with up to $25,000$ points, we implemented a probabilistic selection of inducing points based on [uhrenholt2021] to effectively lower the problem dimension.
Lastly, the solutions found by a variant of CFDNNetAdapt with non-DNN surrogate models are depicted in green. The surrogate models used were based on Gaussian process regression (GPR) [jones1998]. The choice of GPR was motivated by its frequent use (mainly as Kriging models) in SAOAs, as described in, e.g. [morita2022, mastrippolito2021].
Specifically, the sparse variant of the GPR with up to $50$ inducing points was used. In each iteration, four GPR models with different kernels (squared exponential, Matern52, Matern32, and Matern12) were tested. The implementation was done using a Python library gpflow [gpflow2017].
@@
The second indicator was based on the hyper-volume (HV) indicator. In two-dimensional objective space, HV represents an area that is non-dominated by points of a given Pareto-optimal front ($\mathcal{O}^{*}$) with respect to the preset reference point ($\bm{r}$). The formal definition of $\text{HV}(\mathcal{O}^{*},\bm{r})$ is given in [guerreiro2021]. For our comparison, the scaled difference of HV to optimal HV ($\Delta\text{HV}$) was examined. It was computed as
@@
where $\mathcal{O}^{*}_{\mathrm{opt}}$ comprises the optimal objective values. The reference point $\bm{r}$ is problem dependent and was set as the maximal cost function values encountered by NSGA-II. Lastly, the term $\text{HV}(\min(\mathcal{O}^{*}_{\mathrm{opt}}),\bm{r})$ represents the area of a rectangle defined by the minimal optimal values of the cost functions and the reference point.
@@
In ZDT3, the SAOAs still converged faster, but the problems were larger. Surrogate models based on GPRs and RBFs had considerable problems with the discontinuity of the Pareto-optimal front. CFDNNetAdapt with DNNs appears to be the most flexible. In ZDT4, all SAOAs failed. None of the surrogate models was able to give accurate enough predictions and most of the solutions predicted to lie on the Pareto-optimal front (empty circles in Figure 3) were actually far off the true values (full circles in Figure 3).
@@
Still, the standardized tests performed suggest that, similarly to other SAOAs, the CFDNNetAdapt algorithm is not universally usable. The main strength of CFDNNetAdapt lies in accelerating optimizations with computationally intensive evaluation of cost functions. Based on examining Figures 3 and 4, a best-practice approach is proposed. First, CFDNNetAdapt is the most advantageous for optimizations in topologically simple parameter and objective spaces (similar to ZDT1 and ZDT2). Second, for cases with complex objective space but simple parameter space (ZDT3), it can still provide satisfactory solutions with significantly fewer data points than NSGA-II. Third, for the cases of topologically complex parameter space (ZDT4 and ZDT6), the algorithm will fail. This failure is to be expected because in such cases, small variances in cost function parameters significantly affect the returned objectives, which makes construction of any data-driven surrogate problematic at best.
In cases where CFDNNetAdapt can be applied, at least $250$ data points should be available before initial training of the surrogate is attempted. Then, the algorithm usually provides the best trade-off between the quality of approximation of the Pareto-optimal solution to the original problem and the cost function evaluations for less than $2,500$ data points. Using more than $2,500$ data points is beneficial to ensure the CFDNNetAdapt convergence. At $10,000$ data points, CFDNNetAdapt is usually outperformed by pure NSGA-II.
@@ 5.7 Comparison of CFDNNetAdapt and the baseline framework
In each of its iterations, CFDNNetAdapt took $n_{\mathrm{smp}}=600$ points from the common data set. The points were taken in the same order as they were generated by the baseline framework. Regarding the CFDNNetAdapt hyperparameters, the total number of iterations was $k_{\mathrm{max}}=12$, $n_{\mathrm{dnn}}=10$ MLPs were created in each iteration, the number of hidden layers of each MLP was fixed to $n_{\mathrm{hlrs}}=3$ and each MLP was used as a surrogate in an optimization run by NSGA-II with $n_{\mathrm{pop}}=500$ and $n_{\mathrm{gen}}=90$. More details about the hyperparameters are given in C.
@@ 5.8 Validation of selected designs
To support the results presented in Figures 6 and 7, the performance of two designs found by the CFD-driven framework and two found by CFDNNetAdapt was subjected to experimental validation. The designs that were chosen are depicted in Figure 8 in blue and green (baseline framework) and orange and red (CFDNNetAdapt).
@@
With each design, measurements were made for several different primary fluid flow rates $\dot{Q}_{\mathrm{inlet}}$. In addition, a corresponding CFD model was constructed for each measured data point. The complete results of the validation are presented in D.
@@ 6 Summary and conclusion
In this paper, we presented CFDNNetAdapt, an adaptive optimization methodology that integrates computational fluid dynamics (CFD) with multi-objective evolutionary algorithms (MOEA) and is accelerated by deep neural networks (DNNs). The methodology represents a surrogate-assisted approach to reducing computational time and enhancing the efficiency of CFD-based multi-objective optimization. CFDNNetAdapt was tested on several ZDT benchmark functions and applied to a real-life problem of shape optimization of a single-phase ejector, the results of which were validated using experimental data.
The tests on ZDT functions showed that CFDNNetAdapt is at the level of or outperforming similar state-of-the-art methods as SOCEMO [muller2017], or algorithms based on Gaussian process regression [morita2022, mastrippolito2021]. Furthermore, it was listed that for sufficiently simple optimization parameter and objective space, CFDNNetAdapt should be initialized with at least $250$ data points available, but it can be expected to provide a better accuracy-to-cost ratio than NSGA-II for up to $10,000$ data points.
@@
Future work will aim to address weak points of the approach, such as the dependency of CFDNNetAdapt on hyperparameter tuning. Additionally, the scalability of CFDNNetAdapt to other more complex real-life geometries was not fully investigated. For complex geometries, it is favorable to utilize CFD models based on an immersed boundary method for easier geometry change. However, CFDNNetAdapt in its present form is not prepared for that.
CFDNNetAdapt is an open-source project. Its source code and the simulation source codes and settings required to reproduce all the computations of this study are available from https://github.com/techMathGroup/CFDNNetAdapt.
@@ Acknowledgments
This research was co-funded by the European Union under the project Metamaterials for thermally stressed machine components (reg. no. CZ.02.01.01/00/23_020/0008501). The work was financially supported by the institutional support RVO:61388998 and RVO:67985874, and the grant project with No. QL24010110 of the National Agency for Agricultural Research (NAAR). Finally, this work was supported from the grant of Specific university research – A1_FCHI_2025_004 and A1_FCHI_2025_002.
@@ Appendix A Experimental set-up and default geometry
Primary ($\dot{Q}_{\mathrm{inlet}}$) and secondary fluid flow rates ($\dot{Q}_{\mathrm{suction}}$) were measured as the pressure difference of the calibrated orifices; the relative uncertainty of the flow rate is less than 1% of the reading. The pressure in front of the nozzle ($p_{\mathrm{inlet}}$) was measured using the Cressto SR transducer calibrated for the relative pressure range from -100 kPa to 1000 kPa with the uncertainty of $\pm$5 kPa; for the measurement of the pressures in the suction chamber ($p_{\mathrm{suction}}$) and at the diffuser outlet ($p_{\mathrm{outlet}}$), the Cressto SR transducers were calibrated for the relative pressure range from -100 kPa to 300 kPa with the uncertainty of $\pm$0.75 kPa.
@@ B.6 Model validation
Appendix C CFDNNetAdapt settings
hyperpar. value hyperpar. value hyperpar. value
$n_{\mathrm{hlrs}}$ 33 $n_{\mathrm{min}}$ 22 $n_{\mathrm{gen}}$ $250$
$n_{\mathrm{dnn}}$ 44 $n_{\mathrm{max}}$ 2020 $n_{\mathrm{ver}}$ $16$
$\{\bar{n}_{1},\,\bar{n}_{2},\,\bar{n}_{3}\}$ {11, 11, $\{11$,\,11,\,11\} $n_{\mathrm{smp}}$ $1,000$ $k_{\mathrm{max}}$ $25$
$r_{n}$ 44 $n_{\mathrm{pop}}$ $100$ $\epsilon$ $10^{-4}$
Table 1: Values of CFDNNetAdapt parameters used for the ZDT functions.
hyperpar. value hyperpar. value hyperpar. value
$n_{\mathrm{hlrs}}$ 33 $n_{\mathrm{min}}$ 22 $n_{\mathrm{gen}}$ $90$
$n_{\mathrm{dnn}}$ $10$ $n_{\mathrm{max}}$ 2020 $n_{\mathrm{ver}}$ 88
$\{\bar{n}_{1},\,\bar{n}_{2},\,\bar{n}_{3}\}$ {11, 11, $\{11$,\,11,\,11\} $n_{\mathrm{smp}}$ $600$ $k_{\mathrm{max}}$ $12$
$r_{n}$ 44 $n_{\mathrm{pop}}$ $500$ $\epsilon$ $0.05$
The setting of CFDNNetAdapt hyperparameters used for the ZDT functions is listed in Table 1 and for the real-life application in Table 2. In both cases, the process of tweaking the algorithm hyperparameters was similar.
Several parameters were tuned by manually testing a few MLPs with different layer sizes and observing their behavior in optimization. In this way, the hyperparameters $n_{\mathrm{hlrs}}$, $\{\bar{n}_{1},\,\bar{n}_{2},\,\bar{n}_{3}\}$, $r_{n}$, $n_{\mathrm{min}}$, $n_{\mathrm{max}}$, $n_{\mathrm{pop}}$ and $n_{\mathrm{gen}}$ were set.
Then, we ran CFDNNetAdapt without the verification step to tune the hyperparameters $n_{\mathrm{dnn}}$, $n_{\mathrm{smp}}$, $\Delta_{\mathrm{smp}}$ and $\Delta r_{n}$. The results of the last adjustment step were used to perform a manual verification to set the hyperparameters $n_{\mathrm{ver}}$ and $\epsilon$. Lastly, $k_{\mathrm{max}}$ was set to be three times the expected number of steps, since $n_{\mathrm{smp}}$ is limited by the number of available CFD data points.
Appendix D Additional data from validation of optimization results

### v2

@@ 4.2 Estimates of Pareto-optimal solutions
Figure 3: Comparison of Pareto-optimal front estimates for the ZDT problems. From left to right, there are optimal solutions (black), solutions found by NSGA-II (red), solutions found by CFDNNetAdapt (blue), solutions found by SOCEMO (orange), and solutions found by CFDGPRAdapt (green). The empty circles depict the Pareto-optimal front as predicted by the surrogate model, the full circles show the corresponding true values.
In Figure 3, for each SAOA, two estimates of Pareto-optimal fronts corresponding to the same Pareto-optimal set are depicted: the one predicted by the best found surrogate model (empty circles) and its counter-part re-evaluated by the true cost functions (full circles). The solutions (Pareto-optimal fronts) found by NSGA-II are shown in red. CFDNNetAdapt solutions are depicted in blue. The solutions of SOCEMO [muller2017] are shown in orange. Lastly, the solutions found by CFDGPRAdapt are depicted in green.
With $25,000$ function evaluations, NSGA-II converged to a good estimate of the Pareto-optimal front for all the tested functions. As a result, high-quality training samples for SAOAs were always available. On the other hand, for ZDT4 and all the $25,000$ samples used for training, none of the SAOAs was able to converge to the Pareto-optimal solution. For the other ZDT functions, CFDNNetAdapt consistently outperformed the other tested SAOAs.
Figure 4: Comparison of Pareto-optimal set estimates for ZDT problems and all the available data used for SAOAs training. Pareto-optimal sets are shown in black. Next, from left to right, there are solutions found by NSGA-II (red), solutions found by CFDNNetAdapt (blue), solutions found by SOCEMO (orange), and solutions found by CFDGPRAdapt (green).
In Figure 4 are shown the estimates of the Pareto-optimal sets corresponding to the results in Figure 3 for which at least one SAOA converged. Here, note that for ZDT1, both CFDNNetAdapt and CFDGPRAdapt were able to estimate the Pareto-optimal set similarly to NSGA-II while SOCEMO was not able to identify the solutions with $p_{0}\gtrsim 0.8$. On the other hand, CFDNNetAdapt solutions for ZDT2 are plagued by a number of parasitic points. Still, for ZDT3, CFDNNetAdapt was the only SAOA able to correctly identify individual groups of Pareto-optimal solutions.
Similar results to that shown in Figures 3 and 4 but for the LZ functions are provided in Supplementary material. The results for LZ functions confirm the capability of CFDNNetAdapt to identify even complex-shaped Pareto-optimal sets; see Figures 7 and 8 in Supplementary material.
4.3 Evaluating algorithm convergence
@@
The second indicator was based on the hyper-volume (HV) indicator. In two-dimensional objective space, HV represents an area that is non-dominated by points of a given Pareto-optimal front ($\mathcal{O}^{*}$) with respect to a preset reference point ($\bm{r}$). In agreement with [guerreiro2021], the hypervolume indicator can be defined as
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{HV}(\mathcal{O}^{*},\bm{r})=\Lambda\left(\displaystyle\bigcup_{\begin{subarray}{c}\bm{p}\in\mathcal{O}^{*}\\[2.84544pt] {\bm{p}}\text{\ dominates\ }\bm{r}\end{subarray}}\left[\bm{p},\bm{r}\right]\right),\,$} (7)
where $\Lambda(\cdot)$ is the Lebesgue measure and $\left[\bm{p},\bm{r}\right]$ is a box defined by the points $\bm{p}$ and $\bm{r}$. For each optimization problem, the reference point was chosen as $\bm{r}=\left[\displaystyle\max_{\text{NSGA-II}}(f_{1}),\displaystyle\max_{\text{NSGA-II}}(f_{2})\right]$. For a given $\mathcal{O}^{*}$ and $\bm{r}$, the dimension-sweep algorithm [fonseca2006] was used to compute the HV indicator.
For our comparison, the scaled difference of HV to optimal HV ($\Delta\text{HV}$) was examined. It was computed as
@@
where $\mathcal{O}^{*}_{\mathrm{opt}}$ comprises the optimal objective values, and $\text{HV}(\min(\mathcal{O}^{*}_{\mathrm{opt}}),\bm{r})$ represents the area of a rectangle defined by the minimal optimal values of the cost functions and the reference point.
4.4 Discussion of algorithm performance
@@
In ZDT3, the SAOAs still converged faster, but the problems were larger. Surrogate models based on GPRs and RBFs had considerable problems with the discontinuity of the Pareto-optimal front. CFDNNetAdapt appears to be the most flexible. In ZDT4, all SAOAs failed. None of the surrogate models was able to give accurate enough predictions and most of the solutions predicted to lie on the Pareto-optimal front (empty circles in Figure 3) were actually far off the true values (full circles in Figure 3).
@@
Still, the standardized tests performed suggest that, similarly to other SAOAs, the CFDNNetAdapt algorithm is not universally usable. The main strength of CFDNNetAdapt lies in accelerating optimizations with computationally intensive evaluation of cost functions. Based on examining Figures 3 and 5, a best-practice approach is proposed. First, CFDNNetAdapt is the most advantageous for optimizations in topologically simple parameter and objective spaces (similar to ZDT1 and ZDT2). Second, for cases with a complex objective space but a simple parameter space (ZDT3), it can still provide satisfactory solutions with significantly fewer data points than NSGA-II. Third, for the cases of topologically complex parameter space (ZDT4 and ZDT6), the algorithm will fail. This failure is to be expected because in such cases, small variances in cost function parameters significantly affect the returned objectives, which makes the construction of any data-driven surrogate problematic at best.
[…]
CFDNNetAdapt 9.03E-04 (9.99E-05) 2.42E-04 (2.26E-05) [14; 13,778]
ZDT1 SOCEMO 2.18E-02 (9.71E-03) 1.78E-03 (8.61E-04) [1,333; 4,443]
CFDGPRAdapt 1.14E-03 (1.24E-04) 2.91E-04 (4.50E-05) [54; 12,445]
CFDNNetAdapt 9.10E-04 (9.82E-05) 2.05E-04 (2.34E-05) [23; 13,334]
ZDT2 SOCEMO 2.44E-02 (6.90E-02) 3.05E-03 (8.42E-03) [1,333; 9,331]
CFDGPRAdapt 3.85E-02 (9.07E-02) 4.83E-03 (1.09E-02) [89; 1,823]
CFDNNetAdapt 7.84E-04 (1.96E-04) 2.71E-04 (9.54E-05) [14; 11,556]
ZDT3 SOCEMO 1.41E-02 (3.95E-03) 6.09E-03 (1.63E-03) [1,333; 4,443]
CFDGPRAdapt 1.34E-03 (3.80E-04) 4.42E-04 (1.24E-04) [27; 9,778]
CFDNNetAdapt 2.60E+00 (1.55E+00) 4.38E-04 (1.46E-04) [89; 667]
ZDT4 SOCEMO 8.35E-01 (1.06E+00) 9.66E-03 (2.30E-02) [N/A; N/A]
CFDGPRAdapt 2.29E+00 (1.75E+00) 8.22E-04 (8.09E-04) [134; 667]
CFDNNetAdapt 1.22E-01 (2.81E-02) 3.71E-04 (3.28E-04) [14; 178]
ZDT6 SOCEMO 5.36E-01 (2.76E-01) 5.60E-01 (4.02E-01) [N/A; N/A]
CFDGPRAdapt 2.50E-01 (8.49E-02) 1.39E-02 (1.25E-02) [27; 223]
[…]
Figure 6: Comparison of solutions to the ZDT problems and reduced training datasets. Sizes of training datasets are given in parentheses in the first column of the figure. For description of colors and symbols, see Figure 3.
Figure 7: Comparison of Pareto-optimal set estimates for ZDT problems and reduced training datasets. Sizes of training datasets are given in parentheses next to function names. For description of colors and symbols, see Figure 4.
[…]
The estimates of Pareto-optimal fronts and Pareto-optimal sets obtained using this reduced number of tranining samples are shown in Figures 6 and 7. For ZDT1 and 250 training samples, both CFDNNetAdapt and SOCEMO were able to provide acceptable estimates of Pareto-optimal set while incorrectly placing the Pareto-optimal front. Such a situation occurs when the surrogate model can correctly estimate the trends in the mapping between parameter and objective space but the mapping itself is burdened by an error. The third algorithm, CFDGPRAdapt, provided the worst estimate of the Pareto-optimal front and also its estimate of the Pareto-optimal set is of lower quality compared to both CFDNNetAdapt and SOCEMO. However, comparing the number of required function evaluations and the quality of the $\mathcal{O}^{*}$ and $\mathcal{P}^{*}$ estimates, all the three SAOAs significantly outperformed pure NSGA-II.
For ZDT2, the situation was similar to the one observed for ZDT1. Still, this time only CFDNNetAdapt was able to provide a suitable estimate of $\mathcal{P}^{*}$. For ZDT3, a significantly higher number of function evaluations was required for any of the SAOAs to be able to provide an acceptable $\mathcal{O}^{*}$ and $\mathcal{P}^{*}$ estimates. Furthermore, in this case only CFDNNetAdapt was able to outperform NSGA-II. Specifically, the provided estimates of $\mathcal{O}^{*}$ shown in Figure 6 were similar to each other. However, as can be seen in the last row of Figure 7, CFDNNetAdapt was able to provide a better estimate of $\mathcal{P}^{*}$ than NSGA-II.
To conclude, in the cases where surrogate-aided optimization can be applied, CFDNNetAdapt usually provides better trade-off between the quality of approximation of the Pareto-optimal solution and the cost function evaluations than all the other tested approaches. In particular, for ZDT functions and approximately $100$ data points available for training, it significantly outperforms NSGA-II and this effectivity is maintained up to $\approx 2,500$ datapoints. Then, using more than $2,500$ data points is beneficial to ensure the CFDNNetAdapt convergence but the advantages over NSGA-II are less apparent. Finally, at $\approx 10,000$ data points, CFDNNetAdapt is usually outperformed by pure NSGA-II. Still, the ranges of required cost function evaluations for which CFDNNetAdapt outperforms NSGA-II are consistently wider than for both SOCEMO and CFDGPRAdapt. The superior performance of CFDNNetAdapt compared to SOCEMO and CFDGPRAdapt was observed not only for ZDT functions, but also for LZ functions. A detailed analysis for LZ functions is available in Supplementary material.
@@ 5.7 Comparison of CFDNNetAdapt and the baseline framework
hyperpar. value hyperpar. value hyperpar. value
$n_{\mathrm{hlrs}}$ 33 $n_{\mathrm{min}}$ 22 $n_{\mathrm{gen}}$ $90$
$n_{\mathrm{dnn}}$ $10$ $n_{\mathrm{max}}$ 2020 $n_{\mathrm{ver}}$ 88
$\{\bar{n}^{0}_{1},\,\bar{n}^{0}_{2},\,\bar{n}^{0}_{3}\}$ {11, 11, $\{11$,\,11,\,11\} $n_{\mathrm{smp}}$ $600$ $k_{\mathrm{max}}$ $12$
$r_{n}$ 44 $n_{\mathrm{pop}}$ $500$ $\epsilon$ $0.05$
In each of its iterations, CFDNNetAdapt took $n_{\mathrm{smp}}=600$ points from the common data set. The points were taken in the same order as they were generated by the baseline framework. Regarding the CFDNNetAdapt hyperparameters, they are listed in Table 5. More details on the hyperparameters selection are given in Table 1 and its description.
@@ 5.8 Validation of selected designs
To support the results presented in Figures 9 and 10, the performance of two designs found by the CFD-driven framework and two found by CFDNNetAdapt was subjected to experimental validation. The designs that were chosen are depicted in Figure 11 in blue and green (baseline framework) and orange and red (CFDNNetAdapt).
@@
With each design, measurements were made for several different primary fluid flow rates $\dot{Q}_{\mathrm{inlet}}$. In addition, a corresponding CFD model was constructed for each measured data point. The complete results of the validation are presented in C.
@@ 6 Summary and conclusion
In this paper, we presented CFDNNetAdapt, an adaptive optimization methodology that integrates computational fluid dynamics (CFD) with multi-objective evolutionary algorithms (MOEA) and is accelerated by deep neural networks (DNNs). The methodology represents a surrogate-assisted approach to reducing computational time and enhancing the efficiency of CFD-based multi-objective optimization. CFDNNetAdapt was tested on several ZDT and LZ benchmark functions and applied to a real-life problem of shape optimization of a single-phase ejector, the results of which were validated using experimental data.
The tests on ZDT and LZ functions showed that CFDNNetAdapt is at the level of or outperforming similar state-of-the-art methods as SOCEMO [muller2017], or algorithms based on Gaussian process regression [morita2022, mastrippolito2021]. Furthermore, it was listed that for sufficiently simple optimization parameter and objective space, CFDNNetAdapt should be initialized with at least $250$ data points available, but it can be expected to provide a better accuracy-to-cost ratio than NSGA-II for up to $10,000$ data points.
@@
Future work will aim to address weak points of the approach, such as the dependency of CFDNNetAdapt on hyperparameter tuning. Additionally, the scalability of CFDNNetAdapt to other more complex real-life geometries was not fully investigated. For complex geometries, it is favorable to utilize CFD models based on an immersed boundary method for easier geometry change. However, CFDNNetAdapt in its present form is not prepared for that. Nonetheless, the current implementation of CFDNNetAdapt, which is open-source and available from https://github.com/techMathGroup/CFDNNetAdapt, was shown to significantly accelerate multi-objective optimization involving computationally expensive cost functions.
@@ Acknowledgments
This research was co-funded by the European Union under the project Metamaterials for thermally stressed machine components (reg. no. CZ.02.01.01/00/23_020/0008501). The work was financially supported by the institutional support RVO:61388998 and RVO:67985874, and the grant project with No. QL24010110 of the National Agency for Agricultural Research (NAAR). Finally, this work was supported by the Strategy AV21 programme AI: Artificial Intelligence for Science and Society, and from the grant of Specific university research – A1_FCHI_2025_004 and A1_FCHI_2025_002.
@@ Appendix A Experimental set-up and default geometry
Primary ($\dot{Q}_{\mathrm{inlet}}$) and secondary fluid flow rates ($\dot{Q}_{\mathrm{suction}}$) were measured as the pressure difference of the calibrated orifices; the relative uncertainty of the flow rate is less than 1% of the reading. The pressure in front of the nozzle ($p_{\mathrm{inlet}}$) was measured using the Cressto SR transducer calibrated for the relative pressure range from -100 kPa to $1,000$ kPa with the uncertainty of $\pm$5 kPa; for the measurement of the pressures in the suction chamber ($p_{\mathrm{suction}}$) and at the diffuser outlet ($p_{\mathrm{outlet}}$), the Cressto SR transducers were calibrated for the relative pressure range from -100 kPa to 300 kPa with the uncertainty of $\pm$0.75 kPa.
@@ B.6 Model validation
Appendix C Additional data from validation of optimization results